OpenStudy (anonymous):

what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31

OpenStudy (amistre64):

it might be simpler to find the x value and work it from that angle since the ys cancel right away

OpenStudy (amistre64):

3x - 5y = -16 2x + 5y = 31 ------------- 5x = 15

OpenStudy (anonymous):

I know, x = 3

OpenStudy (amistre64):

then if x = 3 2x + 5y = 31 2(3) + 5y = 31 6 + 5y = 31 5y = 31-6 = 25